OpenStudy (anonymous):

Given that cos θ = frac{1}{9}, calculate the value of sin θ.

OpenStudy (anonymous):

I am really stuck on this problem.

OpenStudy (anonymous):

so to solve sin, first you have to solve for cos

OpenStudy (anonymous):

there are 2 ways to do this, pythagorean or using calculators to find degrees.

OpenStudy (anonymous):

I'll show you both, choose one you like more?

OpenStudy (anonymous):

both

OpenStudy (solomonzelman):

1² + b² = 9² sin = b / 9

OpenStudy (amistre64):

sin^2 + cos^2 = 1 :)

OpenStudy (anonymous):

cos theta= 1/9 cos^-1 (1/9) = theta theta = 83.62 degrees sin 83.62= 0.9938

OpenStudy (anonymous):

and the second was solomon's way :D

OpenStudy (anonymous):

why do you square -1?

OpenStudy (anonymous):

a^2 + b^2 = c^2 pythagorean theorem

OpenStudy (anonymous):

so you do pythagorean theorem then sin

OpenStudy (anonymous):

on 2nd though pythagorean would not be applicable for every problems since we don't know if this is a right triangle or not.

OpenStudy (anonymous):

thought* just use your calc to find cos and sin instead. or you can use sin^2 + cos^2 = 1 , that one is always true

OpenStudy (anonymous):

ok thanks